OpenStudy (anonymous):

Given F=9/5C+32. Solve for C.

OpenStudy (anonymous):

a) subtract 32 b) multiply by \(\frac{5}{9}\)

OpenStudy (anonymous):

why would you multiply by 5/9 ?

OpenStudy (anonymous):

because when you subtract \(32\) you get \[F-32=\frac{9}{5}C\] and you don't want \(\frac{9}{5}C\) you want \(C\)

OpenStudy (anonymous):

so if you multiply by the reciprocal of \(\frac{9}{5}\) namely \(\frac{5}{9}\) you get \(C\)

OpenStudy (anonymous):

so in order to get rid of the fraction 9/5 i would have to multiply by the reciprocal?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because \[\frac{5}{9}\times \frac{9}{5}=1\]

OpenStudy (midhun.madhu1987):

 1. Subtract 32 from both sides 2. Multiply both sides by 5. 3. Divide both sides by 9. and you get it.

OpenStudy (anonymous):